The population of a certain country has grown at a rate proportional to the number of people in the country. At present, the country has 80 million inhabitants. Ten years ago it had 70 million. Assuming that this trend continues:

a. Find an expression for the approximate number of people living in the country at any time t (taking t=0 to be the present time)
b. Find the approximate number of people who will inhabit the country at the end of the next ten-year period.
1
Expert's answer
2020-11-30T09:33:59-0500

a.

Let "P" denotes the population at time "t" years and "P_0" be the population at time "t=0". Then,


"\frac{dP}{dt}\varpropto P \implies \frac{dP}{dt}=kP\implies \frac{dP}{P}=kdt." where "k" is a constant.


Integrating both sides, we have;


"\int\frac{dP}{P}= \int kdt\implies \ln P= kt+c\implies P= e^{kt+c}\implies P= Ae^{kt}" .


At time "t=0, P=P_0"


"\implies P_0=Ae^{k.0}\implies A=P_0\implies P=P_0e^{kt}"


At time "t=-10, P= 70,000,000"


"\implies 7\times10^7=8 \times 10^7e^{-10k}\implies \frac{7}{8}=e^{-10k}\\\implies k= \frac{\ln \frac{7}{8}}{-10}"


At any time "t," the population of the country will be "P=P_0e^{kt}" where "P_0= 8 \times 10^7" and "k= \frac{\ln \frac{7}{8}}{-10}" .


b.


At the end of the next 10 years, "t=10," the population will be approximately;


"P=P_0e^{10k}\implies P= 8 \times 10^7 \times \frac{8}{7}\\\implies P=91,428,571.43 \approx9.143 \times 10^7."